# Complete NLO BFKL impact factors for quarkonium hadroproduction in NRQCD: the case of $^{1}S_{0}^{[1]}$, $^{1}S_{0}^{[8]}$, and $^{3}S_{1}^{[8]}$ states

### v2

@@ 5.4 BFKL counterterm for the real-emission part
$V_{g\text{, MRK}}^{(g,1)}=\frac{2z|{\bm{p}}_{T}|\mu^{2\epsilon}}{(2\pi)^{1-2\epsilon}}h_{m}^{(0)}({\bm{p}}_{T}^{2})\int d^{2-2\epsilon}{\bm{k}}_{T}\,{\cal J}_{RR}(k)\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T}-{\bm{k}}_{T}).$ (58)
@@
$\hskip-14.22636ptV_{g,\text{BFKL-R-2}}^{(m,1)}=\frac{4C_{A}|{\bm{p}}_{T}|\mu^{2\epsilon}}{(2\pi)^{1-2\epsilon}}h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})\delta(1-z)\int\frac{d^{2-2\epsilon}{\bm{k}}_{T}\,{\bm{p}}_{T}^{2}}{{\bm{k}}_{T}^{2}({\bm{k}}_{T}^{2}+{\bm{p}}_{T}^{2})}\frac{1}{2}\ln\frac{rs_{0}}{q_{+}^{2}}$ (69)
$=C_{\overline{MS}}|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})\delta(1-z)\left(\frac{{\bm{p}}_{T}^{2}}{\mu^{2}}\right)^{-\epsilon}\bigg(\uwave{-\frac{C_{A}}{\epsilon}\ln\frac{rs_{0}}{q_{+}^{2}}}+O(\epsilon)\bigg)\delta(1-z).$
@@ 5.5 Collecting the contributions together for the $Q\bar{Q}[{}^{1}S_{0}^{[1]}$] state
The final result for the impact-factor in the gluon case is a sum of the contribution of the subtraction term (44)+(50), renormalised virtual contribution (5.2) and the rapidity-factorisation scheme-transition contribution for the BFKL scheme (70). Below we write it in a form, which emphasises the origin and cancellation of various terms:
$V_{g}^{({}^{1}S_{0}^{[1]},1)}({\bm{q}}_{T},{\bm{p}}_{T},z)=V_{g\text{, ST-1}}^{({}^{1}S_{0}^{[1]},1)}+V_{g\text{, ST-2}}^{({}^{1}S_{0}^{[1]},1)}+V_{g\text{, VR}}^{({}^{1}S_{0}^{[1]},1)}+V_{g,\text{BFKL}}^{({}^{1}S_{0}^{[1]},1)}$
@@
$\times K_{\text{BFKL}}({\bm{q}}_{T},{\bm{k}}_{T},{\bm{p}}_{T})+\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})C_{\overline{MS}}\left(\frac{\mu^{2}}{{\bm{p}}_{T}^{2}}\right)^{\epsilon}\bigg[\underbrace{-\ln\frac{\mu_{F}^{2}}{{\bm{p}}_{T}^{2}}P_{gg}(z)}_{\text{(ST-2)+(DGLAP-CT)}}$
@@ 6.2 Extracting the additional soft divergence from the IFs in the CO case
$\Delta V_{g\text{, ST-2}}^{(m,1,{\bf 8})}=\frac{2|{\bm{p}}_{T}|\mu^{2\epsilon}}{(2\pi)^{1-2\epsilon}}h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})$ (79)
@@
$\Delta V_{g\text{, ST-2}}^{(m,1,{\bf 8})}({\bm{q}}_{T},{\bm{p}}_{T},z)=\frac{2|{\bm{p}}_{T}|\mu^{2\epsilon}}{(2\pi)^{1-2\epsilon}}h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})$ (80)
@@
$\Delta V_{g\text{, ST-2-D}}^{(m,1,{\bf 8})}=2C_{A}\frac{2|{\bm{p}}_{T}|\mu^{2\epsilon}}{(2\pi)^{1-2\epsilon}}h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})\bigg(-\frac{\delta(1-z)}{2\epsilon}+\frac{1}{(1-z)_{+}}\bigg)$
@@
$\Delta V_{g\text{, ST-2-D}}^{(m,1,{\bf 8})}=2C_{A}|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})\bigg(-\frac{\delta(1-z)}{2\epsilon}+\frac{1}{(1-z)_{+}}\bigg)$
@@
$\Delta V_{g\text{, ST-2-D}}^{(m,1,{\bf 8})}=|{\bm{p}}_{T}|h_{m}^{(1)}({\bm{p}}_{T}^{2})\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})C_{\overline{MS}}\left(\frac{\mu^{2}}{{\bm{p}}_{T}^{2}}\right)^{\epsilon}\bigg\{\delta(1-z)\bigg[\uwave{\uwave{\frac{C_{A}}{\epsilon}\bigg(1-\ln\frac{M^{2}+{\bm{p}}_{T}^{2}}{M^{2}}\bigg)}}$
@@ 6.4 Virtual part in the CO case and the full result for the CO IF
$V_{g\text{, VR}}^{(m,{\bf 8})}=\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})\delta(1-z)|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})C_{\overline{MS}}\left(\frac{\mu^{2}}{{\bm{p}}_{T}^{2}}\right)^{\epsilon}\bigg\{-\frac{C_{A}}{\epsilon^{2}}-\frac{C_{A}}{\epsilon}\bigg({\ln\frac{r{\bm{p}}_{T}^{2}}{q_{+}^{2}}}+1$
@@
where the finite parts of the one-loop correction $F_{m}^{\text{(virt.)}}(\tau)$ are given in eqns. (32) and (33) for the CO $S$-wave states. The divergent contribution, highlighted in the second line of eqn. (104), is cancelled by the divergence which appears in eqn. (102) due to the radiation of soft gluons by the CO final state. The structure of other divergences in eqn. (104) is exactly the same as in eqn. (5.2) for the CS case, and therefore the rest of the derivation of the eqn. (24) proceeds in exactly the same way. Hence the sum of eqns. (78), (102) and (6.3) with the additional divergence cancelled constitutes the “correction” term to the IF (4.2), which appears in the CO case:
@@
$+\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T})\delta(1-z)|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})\left(\frac{\mu^{2}}{{\bm{p}}_{T}^{2}}\right)^{\epsilon}\bigg\{-\frac{C_{A}}{\epsilon}\bigg(1-\ln\frac{M^{2}+{\bm{p}}_{T}^{2}}{M^{2}}\bigg)\bigg\}.$ (105)
@@ 7 Conclusion
The results presented here constitute the first next-to-leading-order (NLO) calculation of impact factors (IFs) within the BFKL framework for inclusive production of quarkonia, taking into account the heavy-quark-mass effects exactly. Considering gluon-initiated channels, this is also the first NLO result obtained in any small-$x$ formalism.
@@
A second important application is the computation of cross sections for single forward quarkonium production [^10] At high energies and in the forward region, fixed-order collinear quarkonium results feature perturbative instabilities at NLO Schuler:1994hy ; Mangano:1996kg ; Feng:2015cba ; Lansberg:2020ejc , which are resolved by high-energy resummation Lansberg:2021vie ; Lansberg:2023kzf . at full NLL accuracy, either using unintegrated gluon distributions (see e.g. Anikin:2011sa ; Bolognino:2019pba ) or using High-Energy Factorisation Catani:1990xk ; Catani:1990eg ; Catani:1992rn ; Catani:1994sq ; Collins:1991ty , possibly matched to NLO collinear factorisation computations Petrelli:1997ge using the technique similar to that used in Refs. Lansberg:2021vie ; Lansberg:2023kzf . For this purpose the obtained results should be transferred to the appropriate rapidity-factorisation scheme by the known scheme-transformation terms Nefedov:2024swu .
@@ A.2 Different forms of the LO BFKL kernel
$K^{\text{(dim. reg.)}}_{\text{BFKL}}({\bm{q}}_{T},{\bm{p}}_{T})=\frac{\alpha_{s}C_{A}}{\pi^{2}(2\pi)^{-2\epsilon}}\frac{1}{({\bm{p}}_{T}-{\bm{q}}_{T})^{2}}+2\omega_{g}^{(1)}({\bm{p}}_{T}^{2})\delta^{(2)}({\bm{q}}_{T}-{\bm{p}}_{T}),$ (122)
@@ B.2 Quark channel
We collect the expressions for the coefficients of the 1z1-z-expansion for the $^{3}S_{1}^{[8]}$ squared amplitude (176), in the file MEcoeffs_Rq-3S18g.m as a Mathematica array, the first element of which is the coefficient $w^{({}^{3}S_{1}^{[8]})}_{q,0}$.

### v3

@@ 5.4 BFKL counterterm for the real-emission part
$V_{g\text{, MRK}}^{(g,1)}=\frac{2z|{\bm{p}}_{T}|{\mu^{2\epsilon}}}{(2\pi)^{1-2\epsilon}}h_{m}^{(0)}({\bm{p}}_{T}^{2})\int d^{2-2\epsilon}{\bm{k}}_{T}\,{\cal J}_{RR}(k)\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T}-{\bm{k}}_{T}).$ (58)
@@
$\hskip-14.22636ptV_{g,\text{BFKL-R-2}}^{(m,1)}=\frac{4C_{A}|{\bm{p}}_{T}|{\mu^{2\epsilon}}}{(2\pi)^{1-2\epsilon}}h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T})\delta(1-z)\int\frac{d^{2-2\epsilon}{\bm{k}}_{T}\,{\bm{p}}_{T}^{2}}{{\bm{k}}_{T}^{2}({\bm{k}}_{T}^{2}+{\bm{p}}_{T}^{2})}\frac{1}{2}\ln\frac{rs_{0}}{q_{+}^{2}}$ (69)
$=C_{\overline{MS}}|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T})\delta(1-z)\left(\frac{{\bm{p}}_{T}^{2}}{\mu^{2}}\right)^{-\epsilon}\bigg(\uwave{-\frac{C_{A}}{\epsilon}\ln\frac{rs_{0}}{q_{+}^{2}}}+O(\epsilon)\bigg)\delta(1-z).$
@@ 5.5 Collecting the contributions together for the $Q\bar{Q}[{}^{1}S_{0}^{[1]}$] state
The final result for the impact-factor in the gluon case is a sum of the contribution of the subtraction term (44)+(50), renormalised virtual contribution (55) and the rapidity-factorisation scheme-transition contribution for the BFKL scheme (70). Below we write it in a form, which emphasises the origin and cancellation of various terms:
$V_{g}^{({}^{1}S_{0}^{[1]},1{,\text{analyt.}})}({\bm{q}}_{T},{\bm{p}}_{T},z)=V_{g\text{, ST-1}}^{({}^{1}S_{0}^{[1]},1)}+V_{g\text{, ST-2}}^{({}^{1}S_{0}^{[1]},1)}+V_{g\text{, VR}}^{({}^{1}S_{0}^{[1]},1)}+V_{g,\text{BFKL}}^{({}^{1}S_{0}^{[1]},1)}$
@@
$\times K_{\text{BFKL}}({\bm{q}}_{T},{\bm{k}}_{T},{\bm{p}}_{T})+\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T})C_{\overline{MS}}\left(\frac{\mu^{2}}{{\bm{p}}_{T}^{2}}\right)^{\epsilon}\bigg[\underbrace{-\ln\frac{\mu_{F}^{2}}{{\bm{p}}_{T}^{2}}P_{gg}(z)}_{\text{(ST-2)+(DGLAP-CT)}}$
@@ 6.2 Extracting the additional soft divergence from the IFs in the CO case
$\Delta V_{g\text{, ST-2}}^{(m,1,{\bf 8})}=\frac{2|{\bm{p}}_{T}|\mu^{2\epsilon}}{(2\pi)^{1-2\epsilon}}h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T})$ (79)
@@
$\Delta V_{g\text{, ST-2}}^{(m,1,{\bf 8})}({\bm{q}}_{T},{\bm{p}}_{T},z)=\frac{2|{\bm{p}}_{T}|\mu^{2\epsilon}}{(2\pi)^{1-2\epsilon}}h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T})$ (80)
@@
$\Delta V_{g\text{, ST-2-D}}^{(m,1,{\bf 8})}=2C_{A}\frac{2|{\bm{p}}_{T}|\mu^{2\epsilon}}{(2\pi)^{1-2\epsilon}}h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T})\bigg(-\frac{\delta(1-z)}{2\epsilon}+\frac{1}{(1-z)_{+}}\bigg)$
@@
$\Delta V_{g\text{, ST-2-D}}^{(m,1,{\bf 8})}=2C_{A}|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T})\bigg(-\frac{\delta(1-z)}{2\epsilon}+\frac{1}{(1-z)_{+}}\bigg)$
@@
$\Delta V_{g\text{, ST-2-D}}^{(m,1,{\bf 8})}=|{\bm{p}}_{T}|h_{m}^{(1)}({\bm{p}}_{T}^{2})\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T})C_{\overline{MS}}\left(\frac{\mu^{2}}{{\bm{p}}_{T}^{2}}\right)^{\epsilon}\bigg\{\delta(1-z)\bigg[\uwave{\uwave{\frac{C_{A}}{\epsilon}\bigg(1-\ln\frac{M^{2}+{\bm{p}}_{T}^{2}}{M^{2}}\bigg)}}$
@@ 6.4 Virtual part in the CO case and the full result for the CO IF
$V_{g\text{, VR}}^{(m,{\bf 8})}=\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T})\delta(1-z)|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})C_{\overline{MS}}\left(\frac{\mu^{2}}{{\bm{p}}_{T}^{2}}\right)^{\epsilon}\bigg\{-\frac{C_{A}}{\epsilon^{2}}-\frac{C_{A}}{\epsilon}\bigg({\ln\frac{r{\bm{p}}_{T}^{2}}{q_{+}^{2}}}+1$
@@
where the finite parts of the one-loop correction $F_{m}^{\text{(virt.)}}(\tau)$ are given in eqns. (32) and (33) for the CO $S$-wave states. The divergent contribution, highlighted in the second line of eqn. (104), is cancelled by the divergence which appears in eqn. (102) due to the radiation of soft gluons by the CO final state. The structure of other divergences in eqn. (104) is exactly the same as in eqn. (55) for the CS case, and therefore the rest of the derivation of the eqn. (24) proceeds in exactly the same way. Hence the sum of eqns. (78), (102) and (6.3) with the additional divergence cancelled constitutes the “correction” term to the IF (4.2), which appears in the CO case:
@@
$+\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T})\delta(1-z)|{\bm{p}}_{T}|h_{m}^{(0)}({\bm{p}}_{T}^{2})\left(\frac{\mu^{2}}{{\bm{p}}_{T}^{2}}\right)^{\epsilon}\bigg\{-\frac{C_{A}}{\epsilon}\bigg(1-\ln\frac{M^{2}+{\bm{p}}_{T}^{2}}{M^{2}}\bigg)\bigg\}.$ (105)
@@ 7 Conclusion
The results presented here constitute the first next-to-leading-order (NLO) calculation of impact factors (IFs) within the BFKL framework for inclusive production of quarkonia, taking into account the heavy-quark-mass effects exactly. Considering inclusive quarkonium hadroproduction (i.e., gluon- or quark-initiated channels), this is also the first NLO result obtained in any small-$x$ formalism.
@@
A second important application is the computation of cross sections for single forward quarkonium production [^10] At high energies and in the forward region, fixed-order collinear quarkonium results feature perturbative instabilities at NLO Schuler:1994hy ; Mangano:1996kg ; Feng:2015cba ; Lansberg:2020ejc , which are resolved by high-energy resummation Lansberg:2021vie ; Lansberg:2023kzf . at full NLL accuracy, either using unintegrated gluon distributions (see e.g. Anikin:2011sa ; Bolognino:2019pba ; Bolognino:2021niq; Celiberto:2018muu) or using High-Energy Factorisation Catani:1990xk ; Catani:1990eg ; Catani:1992rn ; Catani:1994sq ; Collins:1991ty , possibly matched to NLO collinear factorisation computations Petrelli:1997ge using the technique similar to that used in Refs. Lansberg:2021vie ; Lansberg:2023kzf . For this purpose the obtained results should be transferred to the appropriate rapidity-factorisation scheme by the known scheme-transformation terms Nefedov:2024swu .
@@ A.2 Different forms of the LO BFKL kernel
$K^{\text{(dim. reg.)}}_{\text{BFKL}}({\bm{q}}_{T},{\bm{p}}_{T})=\frac{\alpha_{s}C_{A}}{\pi^{2}(2\pi)^{-2\epsilon}}\frac{1}{({\bm{p}}_{T}-{\bm{q}}_{T})^{2}}+2\omega_{g}^{(1)}({\bm{p}}_{T}^{2})\delta^{(2{-2\epsilon})}({\bm{q}}_{T}-{\bm{p}}_{T}),$ (122)
@@ B.2 Quark channel
We collect the expressions for the coefficients of the 1z1-z-expansion for the $^{3}S_{1}^{[8]}$ squared amplitude (176), in the file MEcoeffs_Rq-3S18q.m as a Mathematica array, the first element of which is the coefficient $w^{({}^{3}S_{1}^{[8]})}_{q,0}$.